[figure: 9.7 cm]

ELECTRONIC AND TRANSPORT PROPERTIES OF A FERROMAGNET WITH A DOMAIN WALL

V. K. Dugaev^{a}, J. Barnaś^{b}, A. Łusakowski^{c}, Ł. A. Turski^{d} aISEL-DEETC, Rua Cons. Emidio Navarro, 1949-014 Lisbon, Portugal

bPhysics Dept., Adam Mickiewicz University, Umultowska 85, 61-614 Poznań, and Institute of Molecular Physics PAS, Smoluchowskiego 17, 60-179 Poznań, Poland

cInstitute of Physics PAS, al. Lotników 32/46, 02-668 Warsaw, Poland dCenter of Theoretical Physics PAS and College of Science, 02-668 Warsaw, Poland

We present a theoretical description of conduction electrons in a ferromagnetic metal, which interact with a domain wall. The description takes into account Coulomb repulsion between electrons and the exchange interaction. Within the semiclassical approximation we calculate the equilibrium spin and charge distributions and show that the wall gives rise to a nonuniform accumulation of both charge and spin. The calculations are selfconsistent and account for the effects of spin and charge redistribution at the wall on the electronic spectrum of conduction electrons propagating in an effective self-consistent elecrostatic potential and in the exchange field created by a nonuniform magnetization. We also calculate the local charge and spin conductivities, and analyse possible effects associated with the charge and spin accumulation at the wall. The nonuniformity of charge conductivity can give rise to the heating effect (Peltier effect) at the domain wall. It also leads to a nonuniform internal electric field, when an external bias voltage is applied. This, together with an inhomogeneous distribution of charge, can result in a drift of the wall due to electric current.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : Poster

Corresponding author : V. K. Dugaev

Address for correspondence : Dr. Vitalii Dugaev Department of Electronics and Communications Instituto Superior de Engenharia de Lisboa Rua Conselheiro Emidio Navarro 1949-014 Lisbon, Portugal

Email address : vdugaev@deetc.isel.ipl.pt